OpenStudy (anonymous):

When x=12, y+15. If x varies directly with y, what is the sum of x and y when y-x=1?

myininaya (myininaya):

x=ky when x=12, y=15?

OpenStudy (anonymous):

4y=5x y-x=1 (5/4)x-x=1 (1/4)x=1 x=4 so y =5 so x+y=5+4=9

myininaya (myininaya):

i guess you mean when x=12,y=15 12=k*15 k=12/15=4/5 so we have x=4y/5 what is x+y when y-x=1 well x=4y/5 so we have y-4y/5=1 5y/5-4y/5=1 1y/5=1 y=5 x=4(5)/5=4 x+y=4+5=9

OpenStudy (anonymous):

yes, sorry about that. thankyou!